OpenStudy (anonymous):

Given the system of equations: 2x – y = –2 x = 14 + 2y What is the value of the system determinant? What is the value of the y−determinant? What is the value of the x−determinant? What is the solution to the system of equations?

OpenStudy (thecatman):

d ithink

OpenStudy (anonymous):

each one is a equation i know its a lot of work... im asking for the system determent the y determent and the x determent plus the solution so ya

OpenStudy (anonymous):

@sourwing i know how to do the matrix and for the y determent i have -2 and 28 but what do i do know?! whats the actual formula

OpenStudy (anonymous):

@a1234

OpenStudy (anonymous):

@wio i know how to do the matrix and for the y determent i have -2 and 28 but what do i do know?! whats the actual formula

OpenStudy (anonymous):

@csscgirl same

OpenStudy (anonymous):

first, i need a reminder, how do you find the determinant? :P sorry

OpenStudy (anonymous):

lol thats what i am trying to figure out but all i know is you set it up like a matrix and do the cross multiplication but after that i am not sure what to do so with the y determinant i did all the steps and am let with 28 and -2 what do i do with them

OpenStudy (anonymous):

lol... oh... but the solution to the system is (-6,-10)

OpenStudy (anonymous):

i think i might be able to figure it out now, lemme c :P gimme a minute

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

did u use the matrix that is: 2 -1 -2 1 -2 14

OpenStudy (anonymous):

I believe if you have the determinants, then \(y = D_y/D\), where \(D_y\) is the \(y\) determinant and \(D\) is the normal determinant.

OpenStudy (anonymous):

As per Kramer's rule.

OpenStudy (anonymous):

@torchriswil What are you determinants?

OpenStudy (anonymous):

Looks like they are not here...

OpenStudy (anonymous):

oh sorry and yes i was looking for the formulas

OpenStudy (anonymous):

Do you know how to find the determinants or need an explanation?

OpenStudy (anonymous):

explanation or formula i dont even know...

OpenStudy (anonymous):

i need explanation

OpenStudy (anonymous):

Okay, first for the normal determinant. It's simple for a \(2\times 2\) matrix. For larger ones you should look it up yourself. First of all, this is our equation in matrix form.\[ \begin{bmatrix} 2&-1\\ 1&-2 \end{bmatrix} \begin{bmatrix} x\\ y \end{bmatrix} = \begin{bmatrix} -2\\ 14 \end{bmatrix} \]

OpenStudy (anonymous):

Next for the \(2\times 2\) matrix, our determinant is: \[ \begin{vmatrix} a&b\\ c&d \end{vmatrix} = a\cdot d - b\cdot c \]

OpenStudy (anonymous):

thank you, ill look at this tomorrow, but parents are making me go to bed... it already 10!

OpenStudy (anonymous):

Ok, good night then.

OpenStudy (anonymous):

thank you so much goodnight :)

OpenStudy (anonymous):

this makes it eaier but i still dont under stand... there is different determenants for y x and normal correct?

OpenStudy (anonymous):

i mean formulas for the determinants

OpenStudy (anonymous):

For the \(x\) determinant, we swap the \(x\) row with the answer row.\[ \begin{bmatrix} \color{red}2&-1\\ \color{red}1&-2 \end{bmatrix} \begin{bmatrix} x\\ y \end{bmatrix} = \begin{bmatrix} \color{blue}{-2}\\ \color{blue}{14} \end{bmatrix} \]And then find the determinant: \[ D_x= \begin{vmatrix} -2&-1\\ 14&-2 \end{vmatrix} \]

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

yes thank you you explain it way better than my teacher

OpenStudy (anonymous):

Tell me what you get.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i have too go to bed but ill get back to you on the answer tomorrow promise

OpenStudy (anonymous):

Sure thing, bye.

OpenStudy (anonymous):

okay so when you are on can you tell me if i did this part right

OpenStudy (anonymous):

a. 2 -1 -1-4= -5 1 2 is this right @wio well for the first one is it?

OpenStudy (anonymous):

\[ \begin{vmatrix} 2&-1\\ 1&2 \end{vmatrix} = (2)(2)-(1)(-1) = 4+1 = 5 \]

OpenStudy (anonymous):

ya my friend told me that say 6- (-1) would be positve becuase a double negative equals a positive so the answers would be a.5 b.-30 c.-26 did i get them right @wio

OpenStudy (anonymous):

@wio d i corrected it did i get them right